OpenStudy (anonymous):

Help, problem : show that lim x to infinity (1+1/x)^e^x = infinity

OpenStudy (anonymous):

take the log i guess then take the limit

OpenStudy (anonymous):

i.e. take the limit \[\lim_{x\to \infty}e^x\ln(1+\frac{1}{x})\]

OpenStudy (anonymous):

that is in the form \(\infty\times 0\) you can change it to \[\lim_{x\to\infty}\frac{\ln(1+\frac{1}{x})}{e^{-x}}\] and use l'hopital

OpenStudy (anonymous):

What is next after applying l'hopital rule?

myininaya (myininaya):

what did you get after using l'hospital?

OpenStudy (anonymous):

\[\frac{ \frac{ 1 }{ x+1 }- \frac{ 1 }{ x } }{- e ^{-x} }\]

OpenStudy (anonymous):

Then i applied it again and got \[\frac{ \frac{ -1 }{ (x+1)^2 } +\frac{ -1 }{ x^2 }}{ e ^{-x} }\]

myininaya (myininaya):

ok or you could say \[\frac{\frac{1}{x+1}-\frac{1}{x}}{-e^{-x}}=\frac{\frac{\frac{-1}{x^2}}{1+\frac{1}{x}}}{-e^{-x}}\]

myininaya (myininaya):

then make that fraction look a lot nicer

myininaya (myininaya):

\[\frac{\frac{\frac{-1}{x^2} \cdot x^2}{(1+\frac{1}{x})x^2}}{-e^{-x}}\]

myininaya (myininaya):

\[\frac{\frac{-1}{x^2+x}}{-e^{-x}}\]

myininaya (myininaya):

now get rid of that compound fraction then you can think about using l'hospital again

OpenStudy (anonymous):

just keep using l'hospital rule or when will it actually show that the lim =\[\infty \]

myininaya (myininaya):

you can keep using l'hosptial if you keep getting 0/0 or inf/inf

myininaya (myininaya):

you need to clean that one fraction up before applying though

myininaya (myininaya):

you need to clean so you can see if you can actually use l'hospital

myininaya (myininaya):

i helped you already clean up your \[\frac{ \frac{ 1 }{ x+1 }- \frac{ 1 }{ x } }{- e ^{-x} } \] and got \[\frac{\frac{-1}{x^2+x}}{-e^{-x}} \] all you have to do is get rid of the compound fraction and see if we have inf/inf or 0/0 then use l'hospital if we do

myininaya (myininaya):

have you cleaned that fraction up?

OpenStudy (anonymous):

working on it :)

myininaya (myininaya):

i will help just a little more \[\frac{\frac{-1}{x^2+x}}{-e^{-x}} =\frac{-1}{x^2+x} \div (-e^{-x})=\frac{1}{x^2+x} \cdot \frac{1}{e^{-x}}=?\]

OpenStudy (anonymous):

\[\frac{ 1 }{ e ^{-x} (x^2+x)}\]

OpenStudy (anonymous):

right?

myininaya (myininaya):

ok but you could rewrite that as ?

myininaya (myininaya):

like we plug in really big number we have something that looks like this 1/(0*inf) that bottom is indeterminate form so you need to rewrite that one fraction above

OpenStudy (anonymous):

\[\frac{ 1 }{ xe ^{-x} (x+1)}\] would this make any difference?

myininaya (myininaya):

nope

myininaya (myininaya):

1/e^-x=?

OpenStudy (anonymous):

ohhhh bring the e^-x up??

OpenStudy (anonymous):

e^x/(x^2)+2?

myininaya (myininaya):

well e^x/(x^2+x) so what form is this?

OpenStudy (anonymous):

\[\infty/\infty \] ?

myininaya (myininaya):

yep so we can use l'hospital

myininaya (myininaya):

as you will see it the outcome would look nicer than before

myininaya (myininaya):

try it

OpenStudy (anonymous):

got e^x / 2x+1 , then e^x/2 and it stops there right?

myininaya (myininaya):

yep but guess what we have .5*e^x what do you know about the right end behavior of e^x?

myininaya (myininaya):

in other words what does e^x approach as x approaches infinity?

OpenStudy (anonymous):

where did you get .5*e^x? i graphed e^x and it seems like its approaching infinity

myininaya (myininaya):

1/2=.5

myininaya (myininaya):

since e^x approaches infinity as x approaches infinity then ke^x approaches infinity where k>0 is a constant

OpenStudy (anonymous):

That was so difficult, but thank you so much for your help.

myininaya (myininaya):

your welcome

OpenStudy (anonymous):

Glad there are people out there like you willing to spend their time helping out other people!